Exercise 5.16 (Algorithmic)> The following table provides a probability distribution for the random variable y f(y) 0. 20 0. 20 0. 30 0. 30 4 9 a. Compute E(y) (to 1 decimal). b. Compute Vary) and ơ (to 2 decimals). Var(y)
(Exercise 5.49 Airline passengers arrive randomly and independently at the passenger-screening facility at a major international airport. The mean arrival rate is 11 passengers per minute. a. Compute the probability of no arrivals in a one-minute period (to 6 decimals). b. Compute the probability that three or fewer passengers arrive in a one-minute period (to 4 decimals). c. Compute the probability of no arrivals in a 15-second period (to 4 decimals). d. Compute the probability of at least one arrival in a 15-second period (to 4 decimals).
Answer #1

a.

E(Y)=sum x*p(x)

= 3 * 0.20 + 4 * 0.20 + 6 * 0.30 +9 * 0.30

= 5.9

b)

Var ( y ) = E ( Y2 ) - ( E ( y ) )2

  E ( Y2 )= 32* 0.20 + 42 * 0.20 + 62 * 0.30 +92 * 0.30

= 40.1

Var ( y) = 40.1 - ( 5.9)2 = 5.29

σ = V/5.29 2.3